OpenStudy (unklerhaukus):

\[\frac{\mathrm dx}{\mathrm dt} = -4\pi^2x+x^2\] \[x_0 = \pi^2,\qquad v_0=0\]

OpenStudy (empty):

Is \(\pi\) the circle constant?

OpenStudy (unklerhaukus):

yes

OpenStudy (empty):

I guess it appears to be a separable differential equation, unless there's some trick going on haha. This is kind of odd looking.

OpenStudy (unklerhaukus):

numeric methods

OpenStudy (empty):

It has a general solution \[x=\frac{4 \pi^2}{e^{4\pi^2 (c+t)}+1}\]

OpenStudy (astrophysics):

Yeah, but how did you get that, I thought separable as well, but what is the numeric method :O

OpenStudy (empty):

I don't know, there are many numerical methods

OpenStudy (unklerhaukus):

the questing is asking me to use the leapfrog integration method, but i'm not convinced it is stable

OpenStudy (usukidoll):

@Astrophysics there was an x left over... usually when it's separable it would be in the form of h(y) dy = f(x) dx and then integrate both sides but that's not the case in this equation :/

OpenStudy (astrophysics):

Ah ok, and I see leap frog method, is this a mechanics problem..mhm

OpenStudy (empty):

Yeah, that's how I got my answer, partial fractions. This is separable @UsukiDoll

OpenStudy (astrophysics):

Wolfram confirms, I don't really know leapfrog integration very well, this does remind me of a simple harmonic oscillator, @Empty you know how to do it using the method

OpenStudy (empty):

Ok but the real question is, "How do I show that leapfrog integration is stable" and that I don't know.

OpenStudy (usukidoll):

X( completely missed that... yes it is separable... (after hours...mind can't think). But what is leapfrog integration?

OpenStudy (usukidoll):

\[\frac{\mathrm dx}{\mathrm dt} = -4\pi^2x+x^2 \] \[\frac{dx}{-4 \pi^2x+x^2} = dt\] \[\frac{dx}{x(-4 \pi^2+x)} = dt\] and then partial fractions... (not doing the rest)

OpenStudy (astrophysics):

Lol I don't think he's allowed to use that

OpenStudy (usukidoll):

-_- k forget it *tosses it in the trash*

OpenStudy (usukidoll):

was fun when it was legit.

OpenStudy (astrophysics):

Haha have a medal, @Michele_Laino may be able to help you

OpenStudy (unklerhaukus):

Catastrophic, error sorrys

OpenStudy (usukidoll):

nah... could be something I haven't done or idk... it's similar to asking me to do a PDE involving Laplace Equation at after 3 am. My brain can't function at this hour..